# Violation of local reciprocity in charge–orbital interconversion

### v1

@@ Abstract
Figure 4: The pumping-induced charge-current generation efficiency $\zeta$ in the W/Ni bilayers. (a) Dependence of $\zeta$ on the W thickness $t_{\mathrm{W}}$ in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer. (b) Dependence of $\zeta$ on the Ni thickness $t_{\rm Ni}$ in the W(0.6 nm)/Ni($t_{\mathrm{Ni}}$) (blue) and W(10 nm)/Ni($t_{\mathrm{Ni}}$) (red) bilayers.
@@
In Fig. 4(a), we show the dependence of the pumping efficiency $\zeta$ on the W-layer thickness $t_{\mathrm{W}}$. Notably, $\zeta$ remains negative in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer even for $t_{\mathrm{W}}>6$ nm, where the orbital torque efficiency $\xi_{\mathrm{DL}}^{E}$ becomes positive (see Fig. 2(a)). The negative sign of $\zeta$ corresponds to the negative sign of the charge-current generation mechanism in the W layer. This implies that $\zeta$ at $t_{\mathrm{W}}>6$ nm is dominated by the inverse SHE, which is known to be negative in W. Nevertheless, as shown in Fig. 4(b), $\zeta$ exhibits a strong dependence on the Ni-layer thickness $t_{\mathrm{Ni}}$ in the W(10 nm)/Ni($t_{\mathrm{Ni}}$) bilayer. This behavior demonstrates that $\zeta$ cannot be solely attributed to the spin pumping, which should be independent of $t_{\mathrm{Ni}}$, just as the spin torque—the reciprocal process of the spin pumping—is also independent of $t_{\mathrm{Ni}}$. We further confirm that the pumping signal is negligible in Ni single-layer films within this $t_{\mathrm{Ni}}$ range supplementary .
@@
The observed long-range dependence of $\zeta$ on $t_{\mathrm{Ni}}$ is characteristic of the reciprocal process of the long-range orbital torque, indicating that the orbital pumping contributes to the charge current in the W(10 nm)/Ni bilayer. Notably, the decrease in $\zeta$ with $t_{\mathrm{Ni}}$ indicates that the orbital-to-charge conversion has a negative sign. In addition to this negative orbital-to-charge conversion, we also observe a positive orbital-to-charge conversion in the W/Ni bilayer. As shown in Fig. 4(a), $\zeta$ exhibits a positive peak around $t_{\mathrm{W}}=0.5$ nm. This positive component increases with $t_{\mathrm{Ni}}$ as shown in Fig. 4(b), indicating that the signal arises from the orbital-to-charge conversion with a positive sign induced by the orbital pumping. These results demonstrate that two distinct orbital-to-charge conversion mechanisms with opposite signs coexist in the W/Ni bilayer: a positive contribution for $t_{\mathrm{W}}<1$ nm and a negative contribution for $t_{\mathrm{W}}>1$ nm.
@@
The orbital-to-charge conversion with the negative sign can be attributed to the bulk inverse OHE in the W layer, demonstrating a violation of local reciprocity between the direct and inverse OHEs. The nature of the orbital-to-charge conversion can differ between the bulk and interfaces, accounting for the two orbital-to-charge conversion mechanisms with different signs observed in the W/Ni bilayer. Since the interface contribution originates from the anisotropic crystal potential, which differs from the bulk crystal potential go2024local , we assume that the interface contribution is particularly significant at the SiO2/W interface rather than at the metallic W/Ni interface. This interface effect should be especially strong at small $t_{\mathrm{W}}$, suggesting that the charge-current generation with the positive sign at $t_{\mathrm{W}}<1$ nm can be attributed to the orbital-to-charge conversion at the SiO2/W interface. As $t_{\mathrm{W}}$ increases, this surface contribution is suppressed, while bulk effects, including the bulk inverse SHE and OHE, become more pronounced. This indicates that the bulk inverse OHE dominates the observed orbital-to-charge conversion with the negative sign at larger $t_{\mathrm{W}}$. These results demonstrate that while the sign of the bulk direct OHE is positive, as evidenced by the torque measurements, the pumping measurements show that the sign of the bulk inverse OHE is negative. The observation of opposite signs for the bulk direct and inverse OHEs demonstrates that the local reciprocity is violated in the charge-orbital interconversion in the $\alpha$-W.
Table 1: Decomposition of the bulk and surface contributions to the direct and inverse OHEs in W go2024local . The bulk and surface contributions to the direct(inverse) orbital Hall conductivity are represented as $\sigma_{\mathrm{OH}}^{\mathrm{dir(inv),bulk}}$ and $\sigma_{\mathrm{OH}}^{\mathrm{dir(inv),surface}}$, respectively, along with the spin Hall conductivity components represented as $\sigma_{\mathrm{SH}}^{\mathrm{dir(inv),bulk}}$ and $\sigma_{\mathrm{SH}}^{\mathrm{dir(inv),surface}}$.
Orbital/Spin Conversion Bulk Surface
OHE Direct $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}>0$ $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}<0$
Inverse $\sigma_{\mathrm{OH}}^{\mathrm{inv,bulk}}<0$ $\sigma_{\mathrm{OH}}^{\mathrm{inv,surface}}>0$
SHE Direct $\sigma_{\mathrm{SH}}^{\mathrm{dir,bulk}}<0$ $\sigma_{\mathrm{SH}}^{\mathrm{dir,bulk}}\simeq 0$
Inverse $\sigma_{\mathrm{SH}}^{\mathrm{inv,bulk}}<0$ $\sigma_{\mathrm{SH}}^{\mathrm{inv,surface}}\simeq 0$
Our observations are consistent with the theoretical prediction of a violation of local reciprocity between the direct and inverse OHEs in $\alpha$-W go2024local . In the theory, the non-reciprocity in the charge-orbital interconversion arises from the non-conservation of the OAM due to its strong interaction with the lattice via crystal fields. This interaction is accounted for by defining the orbital current as the sum of the conventional orbital current and the torque dipole, the latter being associated with the interaction between the OAM and the crystal lattice. It has been shown that the torque dipole is particularly significant at interfaces, where the anisotropic crystal potential efficiently mediates angular momentum exchange between the electrons and the lattice, strongly influencing the local profile of the charge-orbital interconversion in the W layer. As a result, the local charge currents induced by the orbital current deviate from the expectation of strict reciprocity. Notably, even the signs of the local responses differ between the direct and inverse OHEs, as summarized in Table I. The positive and negative signs of the bulk direct and inverse orbital Hall conductivities, $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}>0$ and $\sigma_{\mathrm{OH}}^{\mathrm{inv,bulk}}<0$, evidenced by the torque and pumping results, are consistent with this prediction. Importantly, these opposite signs do not violate Onsager’s reciprocity relation, as the direct and inverse responses are reciprocal when summed over the entire film thickness go2024local . These behaviors contrasts sharply with the SHE in $\alpha$-W, where the direct and inverse SHEs are largely reciprocal even at a local level go2024local . This indicates that the observed violation of the local reciprocity between the direct and inverse OHEs highlights the distinct nature of the charge-orbital interconversion compared to the charge-spin interconversion. Table I also suggests that the surface direct OHE with a negative sign may contribute to the negative torque for $t_{\mathrm{W}}<4$ nm.
In summary, we have experimentally demonstrated the violation of local reciprocity in the charge-orbital interconversion in the W/Ni bilayers. By measuring the orbital torque and orbital pumping, we found that the sign of the bulk inverse OHE is opposite to that of the bulk direct OHE. This result confirms that the charge-orbital interconversion is locally non-reciprocal, in sharp contrast to the global reciprocity governed by Onsager’s reciprocal relations. The observation of the local reciprocity violation in the direct and inverse OHEs highlights the unique characteristics of charge-orbital coupled transport, providing essential insights for a deeper understanding of orbitronic phenomena.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI (Grant Number: 22H04964, 23K19037), Spintronics Research Network of Japan (Spin-RNJ), and MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) (Grant Number: JPJ011438). Y.M. and D.G. acknowledge support from the EIC Pathfinder OPEN grant 101129641 “OBELIX”.

### v2

@@ Abstract
Figure 4: The pumping-induced charge-current generation efficiency $\zeta$ in the W/Ni bilayers. (a) Dependence of $\zeta$ on the W thickness $t_{\mathrm{W}}$ in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer. The values of $t_{\mathrm{W}}$, determined from the sputtering rate, are as follows: 0, 0.42, and 0.79 nm; in 0.36 nm steps from 1.15 to 2.24 nm; in 0.73 nm steps from 2.97 to 8.06 nm; and 9.15 and 9.88 nm. (b) Dependence of $\zeta$ on the Ni thickness $t_{\rm Ni}$ in the W(0.6 nm)/Ni($t_{\mathrm{Ni}}$) (blue) and W(10 nm)/Ni($t_{\mathrm{Ni}}$) (red) bilayers. The values of $t_{\mathrm{Ni}}$, determined from the sputtering rate, are as follows: for the W(0.6 nm)/Ni($t_{\mathrm{Ni}}$), from 5.00 to 13.30 nm in 2.08 nm steps, and 17.45 nm; for the W(10 nm)/Ni($t_{\mathrm{Ni}}$) structure, 4.68 and 5.72 nm, in 2.08 nm steps from 8.83 to 12.98 nm, and 17.13 nm.
@@
In Fig. 4(a), we show the dependence of the pumping efficiency $\zeta$ on the W-layer thickness $t_{\mathrm{W}}$. Notably, $\zeta$ remains negative in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer even for $t_{\mathrm{W}}>6$ nm, where the orbital torque efficiency $\xi_{\mathrm{DL}}^{E}$ becomes positive (see Fig. 2(a)). The negative sign of $\zeta$ corresponds to the negative sign of the charge-current generation mechanism in the W layer. This suggests that $\zeta$ at $t_{\mathrm{W}}>6$ nm is dominated by the inverse SHE, which is known to be negative in W. Nevertheless, as shown in Fig. 4(b), $\zeta$ exhibits a strong dependence on the Ni-layer thickness $t_{\mathrm{Ni}}$ in the W(10 nm)/Ni($t_{\mathrm{Ni}}$) bilayer. This behavior indicates that $\zeta$ cannot be solely attributed to spin pumping, which should be independent of $t_{\mathrm{Ni}}$, just as $\xi_{\mathrm{DL}}^{E}$ of spin torque—the reciprocal process of spin pumping—is also independent of $t_{\mathrm{Ni}}$. This interpretation is further supported by control experiments on spin-dominated Pt/Ni81Fe19 bilayers, showing that $\zeta$ is independent of the FM thickness supplementary . We also confirmed that the pumping signal is negligible in Ni single-layer films within this $t_{\mathrm{Ni}}$ range supplementary .
@@
The observed long-range dependence of $\zeta$ on $t_{\mathrm{Ni}}$ is characteristic of the reciprocal process of long-range orbital torque, indicating that orbital pumping contributes to the charge current in the W(10 nm)/Ni($t_{\mathrm{Ni}}$) bilayer. Notably, the decrease in $\zeta$ with $t_{\mathrm{Ni}}$ indicates that the orbital-to-charge conversion has a negative sign. In addition to this negative orbital-to-charge conversion, we also observe a positive orbital-to-charge conversion. As shown in Fig. 4(a), $\zeta$ exhibits a positive peak around $t_{\mathrm{W}}=0.5$ nm. This positive component increases with $t_{\mathrm{Ni}}$ as shown in Fig. 4(b), indicating that the signal arises from the orbital-to-charge conversion with the positive sign induced by orbital pumping. These results demonstrate that two distinct orbital-to-charge conversion mechanisms with opposite signs coexist in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer: the positive contribution for $t_{\mathrm{W}}<1$ nm and the negative contribution for $t_{\mathrm{W}}>1$ nm.
@@
The orbital-to-charge conversion with the negative sign for $t_{\mathrm{W}}>1$ nm can be attributed to the bulk inverse OHE in the W layer. The nature of the orbital-to-charge conversion can differ between the bulk and interfaces, accounting for the two orbital-to-charge conversion mechanisms with the opposite signs observed in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer for $t_{\mathrm{W}}>1$ nm and $t_{\mathrm{W}}<1$ nm. Since the interface contribution originates from the anisotropic crystal potential, which differs from the bulk crystal potential go2024local , we assume that the interface contribution is particularly significant at the SiO2/W interface rather than at the metallic W/Ni interface supplementary . This interface effect should be especially strong at small $t_{\mathrm{W}}$, suggesting that the charge-current generation with the positive sign at $t_{\mathrm{W}}<1$ nm can be attributed to the orbital-to-charge conversion at the SiO2/W interface. As $t_{\mathrm{W}}$ increases, this surface contribution is suppressed, while bulk effects, including the bulk inverse SHE and OHE, become more pronounced. This indicates that the bulk inverse OHE dominates the observed orbital-to-charge conversion with the negative sign at larger $t_{\mathrm{W}}$.
The above pumping results indicate that the bulk inverse OHE has the negative sign ($\sigma_{\mathrm{OH}}^{\mathrm{inv,bulk}}<0$), which is opposite to the positive sign of the bulk direct OHE ($\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}>0$), as evidenced by the torque measurements. These opposite signs clearly demonstrate a violation of local reciprocity, as expressed by Eq. (2), in the charge-orbital interconversion in $\alpha$-W. The signs of the bulk direct and inverse OHEs, obtained directly from our thickness-dependent torque and pumping measurements, are consistent with the theoretical prediction summarized in Table I in Appendix A. Table I also suggests that the surface direct OHE with a negative sign may contribute to the negative torque for $t_{\mathrm{W}}<4$ nm, although its sign cannot be determined solely from the experimental data supplementary . Importantly, even though interfacial contributions cannot be fully quantified, the bulk contributions are reliably identified from the thickness dependence of the torque and pumping signals. Therefore, the key conclusion of this work—the direct and inverse OHEs in the bulk exhibit the opposite signs—remains robust. The observed local nonreciprocity in the OHEs contrasts sharply with the SHEs, where the direct and inverse SHEs remain largely reciprocal even at a local level in $\alpha$-W go2024local . This contrast highlights the distinct nature of charge-orbital interconversion compared with charge-spin interconversion.
In summary, we have experimentally demonstrated the violation of local reciprocity in the charge-orbital interconversion in the W/Ni bilayers. By measuring the orbital torque and orbital pumping, we found that the sign of the bulk inverse OHE is opposite to that of the bulk direct OHE. This result confirms that the charge-orbital interconversion is locally nonreciprocal, $\sigma_{\mathrm{OH}}^{\mathrm{dir}}(z)\neq\sigma_{\mathrm{OH}}^{\mathrm{inv}}(z)$, in sharp contrast to the global reciprocity expressed in Eq. (1), which is governed by Onsager’s reciprocal relations. The local violation of reciprocity between the direct and inverse OHEs is also expected in other heavy and light metals supplementary , suggesting that local nonreciprocity could serve as a powerful experimental signature of the OHE across a wide range of materials. Because local nonreciprocity originates from the nonconservation of OAM, extending such studies to systems where the loss of OAM can be controlled through interactions with magnons or phonons could open new directions for manipulating charge-orbital coupled transport. In spintronics, while the direct and inverse SHEs have generally been regarded as reciprocal processes, recent studies have revealed that nonreciprocal charge-spin conversion can emerge in Rashba systems or through spin-vorticity coupling PhysRevLett.122.217701 ; PhysRevB.96.140409 . These findings highlight that investigating reciprocity in charge-spin and charge-orbital interconversions is essential for achieving a unified understanding of angular-momentum transport in solids.
@@ Acknowledgements.
We thank J. Yoon and T. Gao for valuable discussions, and K. Yamanoi and Y. Nozaki for their assistance in sample preparation. This work was supported by JSPS KAKENHI (Grant Numbers: 22H04964, 23K19037, 25K21707), Spintronics Research Network of Japan (Spin-RNJ), and MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) (Grant Number: JPJ011438). Y.M. and D.G. acknowledge support from the EIC Pathfinder OPEN grant 101129641 “OBELIX”.